OpenStudy (anonymous):

What are the solutions of x2 − 3x − 18 = 0?

OpenStudy (anonymous):

can you show howw to do it?

OpenStudy (mertsj):

(x-6)(x+3)=0 x-6=0 or x+3=0 x=6 x=-3

OpenStudy (anonymous):

ahh kkkk ty